# Prompt-guided Disentangled Representation for Action Recognition

### v3

@@ 4.3 Ablation Study
Impact of VGNorm. We design VGNorm to adapt to SSG, which integrate both video and graph structures. We compare it with several widely used normalization methods, and observe that VGNorm achieves the best performance as shown in Table 5. Compared to BatchNorm [18], VGNorm better preserves the spatial structure of graphs while maintaining temporal consistency, leading to a significant improvement of +5.71% mAP. Compared to LayerNorm [1], which normalizes features independently and ignores the graph structure, VGNorm incorporates spatially consistent mean scaling to preserve temporal consistency and capture the heterogeneity across SSGs from different videos, resulting in a +2.98% mAP improvement. Finally, compared to GraphNorm [3], VGNorm enhances temporal consistency, enabling the model to better capture temporal dynamics, which contributes to a +2.49% mAP improvement. Moreover, by jointly modeling both temporal consistency and the heterogeneity across SSGs, VGNorm produces more stable and adaptive representations. This leads to the best performance on $a_{m}$, further demonstrating the effectiveness of our normalization strategy.
@@
LayerNorm [1] 68.12 57.96
BatchNorm [18] 65.39 58.20
GraphNorm [3] 68.61 58.32
@@ 4.4 Further Analysis
Figure 4: t-SNE visualization of disentangled action features across 157 classes on Charades [40]. (a) Training set. (b) Test set.
@@ 4.5 Visualization
We computed the inter-class and intra-class distances after applying t-SNE [32] for dimensionality reduction on Charades [40], with the visualization results shown in Figure 4. Specifically, for each video with $L$ labels, we disentangle $L$ action features, covering a total of 157 action classes. We further calculate the inter-class and intra-class distances on both the training and test sets. For the training set, the intra-class and inter-class distances are 2.08 ± 1.57 and 58.58 ± 27.99, respectively, while for the test set they are 1.79 ± 1.45 and 74.94 ± 36.98. Notably, the intra-class distances of disentangled features are mostly within 1, whereas the inter-class distances remain substantially larger. This demonstrates that our feature disentanglement effectively increases the separability between action categories while enhancing the compactness of features within the same category, thereby providing more discriminative representations for downstream action recognition and understanding. We provide more visualization results in Sec E.
@@ Appendix A Video Encoder
We adopt Dual-AI [15] as our video encoder, which is a two-stream architecture composed of interleaved temporal and spatial Transformers [7]. This design facilitates effective spatio-temporal information exchange through alternate modeling along temporal and spatial dimensions. On top of the original architecture, we further introduce additional spatial-to-spatial and temporal-to-temporal branches to enhance the flow of information across both dimensions. The overall architecture is illustrated in Figure  5. After the video encoder, an Multi-Layer Perceptron (MLP) fuses the two branches, and a Feed Forward Network (FFN) [43] further integrates the features to produce the input $\mathcal{F}_{o}^{f}$ for Video Graph Parsing Neural Network (VGPNN).
Both the Spatial Transformer and the Temporal Transformer adopt the standard self-attention encoder architecture [43] (as shown in Figure 7), with the primary difference lying in the dimension along which the attention is applied in spatial or temporal dimension, as shown in Figure 7.
Figure 5: Our video encoder builds upon Dual-AI [15] by introducing additional spatial-spatial and temporal-temporal streams.
@@ B.1 Formulation
Link. For edge $e_{i}^{j,k}$, we have $N$ triples across frames, denoted as $\{(v_{1}^{j},v_{1}^{k},r_{1}^{j,k}),\dots,\,(v_{N}^{j},v_{N}^{k},r_{N}^{j,k})\}$. To model the transition of object interactions over time, we first extract interaction features from these edge triples. Specifically, we concatenate the components of each edge triple and feed them into an edge function $\mathtt{Edge(\cdot)}$ (implemented using a Multi-Layer Perceptron, MLPs; unless otherwise specified, all functions in subsequent modules are implemented as MLPs) to obtain frame-wise relation features $f^{j,k}=\{f^{j,k}_{1},f^{j,k}_{2},\dots,f^{j,k}_{N}\}$. We then apply a self-attention encoder [43] to model the dependencies and transitions among these interaction features across frames to obtain interaction feature $\mathcal{F}^{j,k}=\{\mathcal{F}^{j,k}_{1},\mathcal{F}_{2}^{j,k},\dots,\mathcal{F}^{j,k}_{N}\}$ as follow:
@@
Compared to previous work Graph Parsing Neural Network (GPNN) [37], VPGNN introduces stronger temporal modeling capabilities by modeling edges over time, rather than just generating aggregation weights for edges. VGPNN updates node features at each layer, while GPNN primarily updates edges and only updates nodes in the final layer. Additionally, VGPNN retains the full SSG across frames, unlike GPNN, which only preserves nodes from a single frame. This results in VGPNN offering better interpretability and generalization.
@@ C.2 Construction
STIGPN [46] 54.1
LaIAR [47] 57.2
@@ D.1 Additional Experiment
We conduct domain shift experiments on the Charades dataset to evaluate the generalization ability of our method. Specifically, we follow the experimental setup of  [47], splitting Charades into five disjoint subsets with different action groups. The details are shown in Table 10.
@@ D.2 More Details
LayerNorm [1] 68.12 65.44
BatchNorm [18] 65.39 63.05
GraphNorm [3] 68.61 65.32

### v4

@@ 4.3 Ablation Study
Impact of VGNorm. We design VGNorm to adapt to SSG, which integrate both video and graph structures. We compare it with several widely used normalization methods, and observe that VGNorm achieves the best performance as shown in Table 5. Compared to BatchNorm [batchnorm], VGNorm better preserves the spatial structure of graphs while maintaining temporal consistency, leading to a significant improvement of +5.71% mAP. Compared to LayerNorm [layernorm], which normalizes features independently and ignores the graph structure, VGNorm incorporates spatially consistent mean scaling to preserve temporal consistency and capture the heterogeneity across SSGs from different videos, resulting in a +2.98% mAP improvement. Finally, compared to GraphNorm [graphnorm], VGNorm enhances temporal consistency, enabling the model to better capture temporal dynamics, which contributes to a +2.49% mAP improvement. Moreover, by jointly modeling both temporal consistency and the heterogeneity across SSGs, VGNorm produces more stable and adaptive representations. This leads to the best performance on $a_{m}$, further demonstrating the effectiveness of our normalization strategy.
@@
LayerNorm [layernorm] 68.12 57.96
BatchNorm [batchnorm] 65.39 58.20
GraphNorm [graphnorm] 68.61 58.32
@@ 4.4 Further Analysis
Figure 4: t-SNE visualization of disentangled action features across 157 classes on Charades [charades]. (a) Training set. (b) Test set.
@@ 4.5 Visualization
We computed the inter-class and intra-class distances after applying t-SNE [tsne] for dimensionality reduction on Charades [charades], with the visualization results shown in Figure 4. Specifically, for each video with $L$ labels, we disentangle $L$ action features, covering a total of 157 action classes. We further calculate the inter-class and intra-class distances on both the training and test sets. For the training set, the intra-class and inter-class distances are 2.08 ± 1.57 and 58.58 ± 27.99, respectively, while for the test set they are 1.79 ± 1.45 and 74.94 ± 36.98. Notably, the intra-class distances of disentangled features are mostly within 1, whereas the inter-class distances remain substantially larger. This demonstrates that our feature disentanglement effectively increases the separability between action categories while enhancing the compactness of features within the same category, thereby providing more discriminative representations for downstream action recognition and understanding. We provide more visualization results in Sec E.
@@ Appendix A Video Encoder
We adopt Dual-AI [han2022dual] as our video encoder, which is a two-stream architecture composed of interleaved temporal and spatial Transformers [vit]. This design facilitates effective spatio-temporal information exchange through alternate modeling along temporal and spatial dimensions. On top of the original architecture, we further introduce additional spatial-to-spatial and temporal-to-temporal branches to enhance the flow of information across both dimensions. The overall architecture is illustrated in Figure  5. After the video encoder, an Multi-Layer Perceptron (MLP) fuses the two branches, and a Feed Forward Network (FFN) [transformer] further integrates the features to produce the input $\mathcal{F}_{o}^{f}$ for Video Graph Parsing Neural Network (VGPNN).
Both the Spatial Transformer and the Temporal Transformer adopt the standard self-attention encoder architecture [transformer] (as shown in Figure 7), with the primary difference lying in the dimension along which the attention is applied in spatial or temporal dimension, as shown in Figure 7.
Figure 5: Our video encoder builds upon Dual-AI [han2022dual] by introducing additional spatial-spatial and temporal-temporal streams.
@@ B.1 Formulation
Link. For edge $e_{i}^{j,k}$, we have $N$ triples across frames, denoted as $\{(v_{1}^{j},v_{1}^{k},r_{1}^{j,k}),\dots,\,(v_{N}^{j},v_{N}^{k},r_{N}^{j,k})\}$. To model the transition of object interactions over time, we first extract interaction features from these edge triples. Specifically, we concatenate the components of each edge triple and feed them into an edge function $\mathtt{Edge(\cdot)}$ (implemented using a Multi-Layer Perceptron, MLPs; unless otherwise specified, all functions in subsequent modules are implemented as MLPs) to obtain frame-wise relation features $f^{j,k}=\{f^{j,k}_{1},f^{j,k}_{2},\dots,f^{j,k}_{N}\}$. We then apply a self-attention encoder [transformer] to model the dependencies and transitions among these interaction features across frames to obtain interaction feature $\mathcal{F}^{j,k}=\{\mathcal{F}^{j,k}_{1},\mathcal{F}_{2}^{j,k},\dots,\mathcal{F}^{j,k}_{N}\}$ as follow:
@@
Compared to previous work Graph Parsing Neural Network (GPNN) [gpnn], VPGNN introduces stronger temporal modeling capabilities by modeling edges over time, rather than just generating aggregation weights for edges. VGPNN updates node features at each layer, while GPNN primarily updates edges and only updates nodes in the final layer. Additionally, VGPNN retains the full SSG across frames, unlike GPNN, which only preserves nodes from a single frame. This results in VGPNN offering better interpretability and generalization.
@@ C.2 Construction
STIGPN [STIGPN] 54.1
LaIAR [lair] 57.2
@@ D.1 Additional Experiment
We conduct domain shift experiments on the Charades dataset to evaluate the generalization ability of our method. Specifically, we follow the experimental setup of  [lair], splitting Charades into five disjoint subsets with different action groups. The details are shown in Table 10.
@@ D.2 More Details
LayerNorm [layernorm] 68.12 65.44
BatchNorm [batchnorm] 65.39 63.05
GraphNorm [graphnorm] 68.61 65.32